ROVIBRATIONAL SPECTROSCOPY OF THE OH-O_{3} AND C_{2}H_{4}-O_{3} COMPLEXES IN ^{4}He NAN-ODROPLETS

\underline{EMMANUEL I. OBI}, and GARY E. DOUBERLY, Department of Chemistry, University of Georgia, Athens, Georgia 30602, USA.

The weakly bound complexes X-O_{3} (X=OH, C_{2}H_{4}) have been isolated in helium nanodroplets and probed with infrared laser spectroscopy. Measurements are carried out in the high frequency OH/CH stretch region (2900-3600 cm^{−1}) and in the OOO symmetric and antisymmetric stretch region near 1050 cm^{−1}. Many of the observed bands exhibit rotational fine structure, allowing for structural assignments via Stark spectroscopy. Rotationally resolved vibrational bands of the OH-O_{3} complex indicate a rather weak interaction, as the orbital angular momentum of the unpaired electron is unquenched and remains strongly coupled to the OH bond axis.